OpenStudy (usukidoll):

This is about the finding the area formula. How do I get my a and b values for this equation: r = 4 +2cospheta

OpenStudy (usukidoll):

area formula is 

OpenStudy (usukidoll):

@experimentX @UnkleRhaukus @Algebraic!

OpenStudy (experimentx):

first of all plot this equation \[ r = 4 +2\cos \theta\]

OpenStudy (usukidoll):

that's a big circle if I remembered correctly

OpenStudy (usukidoll):